1. Calculate the wavelength of a waveform given a period of 17μs.
1
Expert's answer
2021-10-10T15:39:54-0400

"v=n\lambda=\frac{\lambda}{T}"

v=c

"\lambda=c\times T"


"\lambda=3\times10^8\times17\times10^{-6}=5100m"